Okay so in class we are learning to combine like terms.
Equation: 3x+12-10x
I've tried factoring but I'm very confused.

Answer by jim_thompson5910(35256)

3x and -10x are like terms. Combine them to get -7x. Notice how 3 + (-10) = 3 - 10 = -7

So that's why 3x - 10x = -7x

So in the end, the final answer is -7x + 12